# Operating System And Artificial Intelligence: A Systematic Review

### v1

@@ Abstract
In the dynamic landscape of technology, the convergence of Artificial Intelligence (AI) and Operating Systems (OS) has emerged as a pivotal arena for innovation. Our exploration focuses on the symbiotic relationship between AI and OS, emphasizing how AI-driven tools enhance OS performance, security, and efficiency, while OS advancements facilitate more sophisticated AI applications. We delve into various AI techniques employed to optimize OS functionalities, including memory management, process scheduling, and intrusion detection. Simultaneously, we analyze the role of OS in providing essential services and infrastructure that enable effective AI application execution, from resource allocation to data processing. The article also addresses challenges and future directions in this domain, emphasizing the imperative of secure and efficient AI integration within OS frameworks. By examining case studies and recent developments, our review provides a comprehensive overview of the current state of AI-OS integration, underscoring its significance in shaping the next generation of computing technologies. Finally, we explore the promising prospects of Intelligent OSes, considering not only how innovative OS architectures will pave the way for groundbreaking opportunities but also how AI will significantly contribute to advancing these next-generation OSs.
@@ 1 Introduction
Operating systems (OSes) have long been central to computer systems, efficiently managing hardware resources and providing secure environments for application execution. However, the increasing complexity of modern OSes, the rapid diversification of hardware, and the continuous evolution of Artificial Intelligence (AI) present new opportunities to explore AI’s potential throughout the lifespan of OSes, spanning both development and runtime. As the AI wave continues to surge, AI systems grow increasingly massive and complex, necessitating optimization and efficiency enhancements at the lower layers of the stack, particularly within the operating system.
This review delves into the intricate relationship between AI and OS, examining how AI-driven tools enhance OS performance, security, and efficiency. Conversely, we explore how advancements in OS design facilitate the deployment and optimization of AI applications. By scrutinizing various AI techniques employed to augment OS functionalities—such as memory management, process scheduling, and intrusion detection—this review provides a comprehensive overview of the current state of AI-OS integration. Additionally, we discuss the critical role of the OS in providing essential services and infrastructure necessary for effective AI application operation, from resource allocation to data processing. The article also addresses challenges and future research directions in this domain, emphasizing the imperative for secure and efficient integration of AI capabilities within OS frameworks. Through systematic examination of case studies and recent developments, this review underscores the significance of the AI-OS nexus in propelling the next generation of computing technologies.
In this paper, we embark on a comprehensive exploration of the intersection between AI and OS. Our approach involves collecting and analyzing 108 primary studies in the field, aiming to uncover key insights and trends. The rest of this paper is organized as follows. We provide the research background.Then we propose our research questions, report on the paper selection process, and analyze the distribution of research popularity. Collected papers in the fields of AI4OS and OS4AI are classified and summarized in Sections 4, 5 and 6. Furthermore,we discuss how and what novel OS architectures may create opportunities for AI4OS in section 7 and how LLM can create oppotunies for OS in section 8.
Subsequently, in Section 9, we consolidate these insights by outlining prospective opportunities for future work, particularly emphasizing the synergy between Library OS paradigms and LLM-driven approaches in the pursuit of Intelligent OS. This section underscores the immense potential for synergistic collaborations between these technologies, which could pave the way for highly modular, adaptable, and self-aware operating systems. Finally, we bring this comprehensive exploration to a close in Section 10, synthesizing our findings and offering a compelling conclusion.
2 Background
Operating systems have been providing a crucial layer of abstraction between applications and hardware resources. The design of an OS can significantly impact hardware resource management efficiency and, consequently, the performance of all applications running on it. Additionally, the development of OSes requires extensive engineering efforts from experts. And the complexity of modern OSes poses challenges for developers to effectively improve and optimize performance. The scale and dynamics of computing systems further contribute to this complexity.Moreover, the escalating scale and intricacy of AI systems themselves necessitate optimizations from the OS. As the complexity increases and the need for efficient resource management grows, the integration of AI techniques has emerged as a promising avenue for enhancing OS functionality and performance.
In recent years, researchers have explored various dimensions of AI integration in operating systems, leading to notable advancements.
3 Methodology
We present the research questions, introduce the paper selection process, and statistically analyze the selected papers in this section.
3.1 Research questions
The definition of research questions is the core innovation of a secondary study, as they clearly convey the authors’ perspective on the subject under investigation and the study’s goal.We characterize the purpose, the major research topics, and the scope of this paper as follows.
RQ1: What OS sub-domains are researchers inclined to enhance using AI?
RQ2: How can OS be optimized to improve the efficiency of AI systems?
RQ3: How and which AI techniques are used to improve OS?
RQ4: How and what novel OS architectures create opportunities for AI4OS?
RQ5: How LLMs create opportunities for OS?
@@ 3.2 Paper Selection
We identify four AI-related terms and two OS-related terms, then combine these two categories of terms by logical ORs to create nine search strings. After specifying the range the papers were published: 2019–March 2024, we deliver these nine search strings to Google Scholar, to collect related papers in the intersection of AI and OS. The defined search terms are as follows:
AI or Machine Learning or Deep Learning or Large Language Model
@@
We began with an initial pool of 212 papers related to the intersection of AI and Operating Systems (OS). To narrow down our focus, we applied the following inclusion criteria (IC) and exclusion criteria (EC):
IC1: Peer-Reviewed Research Paper: We considered only peer-reviewed research papers.
IC2: Primary Study: We included primary studies that directly addressed AI-OS integration.
IC3: Relevance to AI and OS: Papers needed to discuss AI techniques enhancing OS or propose novel approaches for improving AI systems using OS methods.
IC4: Publication Date: We restricted our selection to papers published within the past 5 years.
@@
EC1: Avoiding Duplicates: Only the most comprehensive or recent version of each study was included.
EC2: Length Constraint: Papers shorter than 6 pages were excluded.
EC3: Language: We considered papers written in English.
As a result, 108 papers met our criteria and are within the scope of our research. Excluded papers merely mentioned the terms but did not delve into the actual intersection of OS and AI.
Figure 1: Year Distribution of Paper Published
3.3 Paper Analysis
3.3.1 The main venues
Figure 1 depicts the distribution of publication years for research 1. Notably, the number of papers published in the AI4OS and OS4AI domains indicates a growing trend, increased attention in this field. Notably, the number of papers published(including pre-print) in the field of integrating LLM and OS indicate a rapidly growing trend, 17 publications in this field since 2023, when powerful LLM emerged.
Journals and conferences are listed in Tables 1 , the selection criterion is how many times they appear in our paper pool. International Conference on Architectural Support for Programming Languages and Operating Systems(ASPLOS) is the most prevalent conference.
TABLE I: Conferences Distribution
Conference Papers
ASPLOS 22
HPCA 11
USENIX ATC 11
OSDI 5
SOSP 2
HotOS 2
APsys 2
SOSP 2
DAC 1
CCS 1
MobiCom 1
ESEC/FSE 1
ACM TURC 1
ISSTA 1
We first divide all the papers into AI4OS(38), OS4AI(53), LLM AS OS(9) and LLM4OS(8).
4 RQ1: What OS sub-domains are researchers inclined to enhance using AI?
Currently, OSes employ global, static policies based on heuristics. However,AI techniques, adaptable to different application behaviors, hold promise for outperforming these traditional policies. This research question explores and quantifies the diverse applications of AI approaches in enhancing or automating various OS tasks.
4.1 AI for OS Tuning
Growing OS complexity challenges configuration and decision-making, prompting the exploration of AI and ML for enhancements in auto-tuning tasks, including scheduling, energy efficiency, and memory management.
Figure 2: General AI for OS Tuning Workflow
Figure 2 showed the general workflow for AI tools to assist OS.
Scheduling
Scheduling, a core OS function, balances fairness, responsiveness, and throughput by efficiently managing resources.[2][16] Advancements through machine learning can refine ticket distribution, enable adaptability, and adjust scheduling variables precisely.
Springborg et al. [86] introduce Chronus, a Python application that collaborates with the Simple Linux Utility for Resource Management (SLURM) scheduler, prevalent in global supercomputers. Chronus executes comprehensive benchmark tests on HPC clusters, varying parameters like core numbers, processor speeds, and hyperthreading. Throughout, it logs performance metrics and energy use for each setup. This data trains a machine learning model to recognize patterns, aiming to forecast the most energy-efficient configuration for specific jobs and systems. Proven through notable energy savings in benchmarks, this approach holds potential for broader application in HPC systems, promoting eco-friendliness and financial prudence without sacrificing computational power.
@@ 4.1 AI for OS Tuning
Chen et al.[3] presents a ML-based resource-aware load balancer for the Linux kernel with a low-overhead method for collecting training data,an ML model based on a multi-layer perceptron that imitates the Linux’s Completely Fair Scheduler (CFS) load balancer based on the collected training data and an in-kernel implementation of inference on the model. The authors argue that CFS approach maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources.Using eBPF for dynamic tracing, an MLP model replicates CFS decisions, with in-kernel inference for real-time balancing, the model achieves high accuracy and small latency increase, demonstrating effective, low-overhead load management.
@@
Goodarzy et al.[14] also questioned CFS in its ability for proper allocation of CPU, memory, I/O, and network bandwidth. In response to this, the authors propose SmartOS, an operating system that automatically learns what tasks the user deems to be most important at that time. ased on the learned user preferences, SmartOS adjusts the allocation of system resources such as CPU, memory, I/O, and network bandwidth. It prioritizes the resources for the tasks that the user is currently focused on. The authors demonstrate an implementation of such a learning-based OS in Linux and present evaluation results showing that a reinforcement learning-based approach can rapidly learn and adjust system resources to meet user demands.
Storage
@@
Storage systems, along with their associated OS components, are engineered to cater to a broad spectrum of applications and fluctuating workloads. The storage elements within the OS incorporate a range of heuristic algorithms, ensuring optimal performance and adaptability across diverse workloads.
@@
Predictable latency can be very useful for data-center systems serving interactive applications such as messaging and search.Cortez et al.[6] introduce LinnOS, a novel approach to managing SSD performance by incorporating a lightweight neural network within the operating system’s core. This neural network enables real-time inference of SSD performance at an extremely fine-grained level, specifically per-I/O operation, without requiring modifications to the underlying hardware or changes to file systems and applications. LinnOS profiles the latency of a large number of I/O operations submitted to the SSD, leveraging this data to train the neural network. By adopting a ”black-box” perspective towards the device, LinnOS learns and infers the SSD’s behavior to increase predictability, thereby empowering applications to anticipate whether their performance expectations will be met.
Wu et al. [85] introduce LearnWD, a novel approach that harnesses the synergies of machine learning and out-of-place updates to effectively mitigate the write disturbance (WD) problem in NVM. At its core, LearnWD employs clustering algorithms to systematically categorize stale data based on their inherent error proneness. Upon receiving a write request, LearnWD meticulously assesses both the aggressiveness of incoming new data and the error vulnerability of the existing stale data. By thoughtfully leveraging this information, LearnWD strategically orchestrates write operations to proactively minimize the incidence of WD errors.

### v2

# Integrating Artificial Intelligence into Operating Systems: A Comprehensive Survey on Techniques, Applications, and Future Directions

@@ Abstract
In the era of the Internet of Everything, operating systems (OSs) face unprecedented challenges posed by an evolving application landscape and increasingly heterogeneous hardware ecosystems. This shift toward increasingly dynamic and unpredictable operational contexts presents significant challenges for both OS developers and users. Against this backdrop, the fusion of Artificial Intelligence (AI) with Operating Systems emerges as a critical frontier for innovation. This survey delves into the intricate interplay between AI and OSs, illustrating how existing OS mechanisms combined with AI significantly elevate the performance, security, and efficiency of modern operating systems. We investigate a range of AI methodologies applied to optimize core OS functionalities and clarify the correlation to related studies. Our analysis touches on the existing hurdles and prospective avenues in this interdisciplinary domain, underscoring the imperative for robust and seamless integration of AI capabilities into OS architectures.
Through an examination of illustrative case studies and cutting-edge developments, we offer a thorough review of the current status of AI-OS integration, accentuating its pivotal role in steering the evolution of advanced computing paradigms. We also envision the promising prospects of Intelligent Operating Systems, debating how groundbreaking OS designs will usher in novel possibilities and highlight the central role that AI will assume in propelling these next-generation systems forward. This forward-thinking outlook illuminates the profound influence of AI on the foundational elements of computing, heralding the advent of a new epoch characterized by intelligent, self-adapting, and highly adaptive software ecosystems.
@@ 1 Introduction
The evolution of operating systems (OS) has been shaped by the continuous interplay between technological advancements and the changing demands of computing environments. From the early days of manually configured systems to today’s sophisticated platforms, OS development has been driven by the need to efficiently manage resources, enhance usability, and adapt to emerging paradigms. Key milestones, such as the introduction of Unix and the rise of graphical user interfaces, laid the foundation for modern operating systems, enabling multitasking, modularity, and accessibility. Over the decades, the diversification of computing devices, from desktops and mobile devices to embedded systems, has further propelled innovations in OS design to address the unique requirements of specialized hardware and applications.
In recent years, the integration of artificial intelligence (AI) into operating systems has emerged as a transformative trend, unlocking new opportunities for intelligent resource management, adaptive performance, and enhanced user experiences. AI encompasses a range of technologies, including expert systems, machine learning (ML), and large language models (LLMs), each contributing distinct capabilities to OS functionality. By embedding AI technologies, operating systems can evolve from static, rule-based platforms into dynamic, intelligent systems capable of optimizing operations, predicting user needs, and addressing the growing complexity of modern computing environments.
The motivation for integrating AI into OS lies in its potential to address critical challenges in areas such as resource allocation, security, and system adaptability. For example, ML algorithms enable OSs to learn from data and adjust configurations dynamically, while LLMs enhance user interactions through natural language processing, enabling intuitive and conversational interfaces. Additionally, the adoption of AI-driven techniques aligns with emerging trends in computing, such as the Internet of Things (IoT), edge computing, and cloud systems, where managing large-scale, interconnected environments requires intelligent and scalable solutions.
This survey explores the intersection of AI and OS, providing a comprehensive review of recent advancements, challenges, and future opportunities. We categorize the integration of AI technologies into two key aspects of OS development: design and operational optimization. Our analysis is based on an extensive review of primary studies, highlighting the impact of AI on OS performance, efficiency, and adaptability. The remainder of this paper is organized as follows: Section 2 outlines the scope and objectives of the survey. Sections 3 and 4 discuss AI-driven methodologies for OS development and operational enhancements. Section 5 examines the integration of LLMs into OS architectures, and section 6 discusses the clarification of different stages of AI-assisted OSs. Section 7 identifies emerging trends and future directions, while Section 8 concludes the paper, summarizing the key insights and contributions.
Here are the highlights of this paper:
  • We identify and discuss the application of ML tools across various subdomains of operating systems.
  • We provide detailed insights into a variety of ML algorithms used in different areas of operating system research, which includes understanding the specific methodologies employed to achieve performance improvements and security enhancements.
  • We explore how LLMs can be utilized within the context of operating systems, suggesting new research avenues and potential improvements that can be realized through their integration.
  • We identify novel operating system architectures that provide opportunities for the development of intelligent operating systems. This includes examining designs that facilitate better AI integration and enable advanced features such as dynamic resource allocation and improved security protocols.
  • We define and outline the different stages of AI integration within operating systems, providing a roadmap for ongoing and future research efforts aimed at enhancing OS functionality through AI.
2 Related Search
Previous research has made significant strides in exploring the integration of AI with OS (?). For instance, (?) examines the application of quantum computing in cybersecurity, highlighting how AI and qubit-based operating systems can enhance security measures. This work points to the potential of quantum-enhanced systems to effectively address complex security challenges. Additionally, (?) investigates the use of object-oriented programming and AI methodologies to structure and optimize operating systems, advocating for more compact, customized systems that meet the diverse and evolving demands of various computing environments. This study underscores the importance of adaptability and efficiency in modern OS design. In (?), the focus is on the role of AI and ML in OS management, presenting a framework for their integration aimed at enhancing resource allocation, security, performance, and system adaptability. This research emphasizes the need for intelligent systems capable of self-optimization and dynamic responses to changing conditions. Furthermore, (?) assesses the impact of various AI techniques on key aspects of distributed operating systems, including resource management, fault tolerance, and security. (?) focuses on the improvements that ML can bring to OS resource management, while (?), (?), and (?) address security aspects of operating systems. Additionally, (?) explores the application of AI methods in software configurations.
While these works have provided valuable insights and advancements, they often concentrate on specific types of operating systems or may now be considered somewhat outdated, particularly given the recent emergence and rapid development of large language models (?). As we move forward, it will be crucial to explore how LLMs and other cutting-edge AI technologies can be integrated into operating systems to drive the next wave of innovation in computing.
3 Research Method
@@ 3.1 Paper Selection
3.1.1 Search string design
We identify five AI-related terms and two OS-related terms, then combine these two categories of terms by logical ORs to create ten search strings.
AI or Machine Learningor Deep Learning or Large Language Model or LLM
@@ 3.1.1 Search string design
3.1.2 Search database
To gather all relevant articles related to our research, we must select suitable sources. For this purpose, we chose databases commonly used by researchers for conducting surveys and reviews. The selected sources are listed below:
• ACM Digital Library (www.acm.org/dl)
• IEEE Xplore (ieeexplore.ieee.org)
• arXiv by Cornell University (www.arxiv.org)
• Google Scholar (www.scholar.google.com)
The search result from different data sources are listed in Table 1, and has a total of 824 papers.
3.1.3 Inclusion Criteria and Exclusion Criteria
Then following inclusion criteria (IC) and exclusion criteria (EC) are conducted:
IC1: The work is publicated after 2019.
IC2: Primary study.
IC3: The work is about integrating AI to assist OS.
@@ 3.1.3 Inclusion Criteria and Exclusion Criteria
EC1: The duplicate version of a selected paper (only the most comprehensive or recent version of each study should be included).
EC2: The length of the paper is less than 6 pages.
EC3: Paper written in a language other than English and Chinese.
EC4: If an article is repeating in different resources, then the older duplicate was excluded.
After applying these ICs and ECs, 88 research papers are selected. Detailed information on the selected papers is provided in B.
[…]
my root size/.style= minimum height=1cm, text width=1cm , my tnode size/.style= minimum height=2cm, text width=1cm
{forest}
[…]
Figure 1: Literature Survey Tree
4 Machine Learning Enhanced Operating Systems
OSs manage a vast number of processes that continuously start and stop. These systems could benefit from analyzing historical execution behavior to optimize resource management, such as intelligently allocating resources based on usage patterns. However, OSs often rely on hard-coded heuristics that lack adaptability to changing usage scenarios (?). This highlights the need for mechanisms that can leverage the accumulated knowledge and patterns of user behavior to determine optimal configurations for operational timings, including thread scheduling, memory paging, and the frequency of buffer cache flushing. In this section, we will review the efforts documented in the literature that detail the application of ML methods to enhance OS functionalities.
4.1 Scheduling
Scheduling, a core mechanism for OSs, balances fairness, responsiveness, and throughput by efficiently managing resources (?, ?). The effectiveness of a scheduler directly impacts overall system performance, influencing factors such as task latency, resource utilization, and user experience. However, in the context of increasingly complex hardware architectures and diverse software workloads, traditional OS schedulers, which rely heavily on static, heuristic-based algorithms, often struggle to adapt to dynamic and heterogeneous environments. These limitations manifest as inefficiencies in handling workload variability, hardware resource contention, and energy optimization. To address these challenges, recent research has explored the integration of ML techniques into OS scheduling. By leveraging data-driven models, ML-based schedulers aim to automate decision-making processes and transcend the constraints of traditional approaches. These adaptive methods enhance scheduling flexibility, improve prediction accuracy for workload behaviors, and optimize resource allocation in real-time, thus addressing the evolving demands of modern computing systems.
Nevertheless, the existing heuristic-based algorithms are difficult to achieve its most efficiency since they tend to focus on locally optimal solutions, often missing the global optimum due to limited exploration and static rules. These algorithms are typically tailored to specific scenarios and struggle to generalize to dynamic or complex scheduling problems with intricate dependencies. Their reliance on predefined rules and parameters can make them inflexible and sensitive to changes, leading to suboptimal performance in varying conditions. Moreover, heuristics lack scalability and fail to guarantee solution quality, making them unsuitable for applications requiring precision or adaptability. As a result, modern approaches like machine learning or metaheuristics are often preferred for more robust and adaptive scheduling solutions. In contrast to conventional heuristic-based approaches, generative models can also be utilized in scheduling to address challenges such as modeling dynamic and complex environments. Their ability to capture underlying patterns and properties from observed data makes them well-suited for simulating and predicting system behaviors. For instance, generative model-based approaches can effectively represent intricate workload patterns or resource contention scenarios, enabling more accurate modeling and decision-making. Additionally, these models excel in handling the complexity of modern computational environments, offering improved adaptability and precision compared to traditional methods.
@@ 4.1 Scheduling
For OS in individual hosts, advancements through ML can refine ticket distribution, enable adaptability, and adjust scheduling variables precisely. Their purpose is to ensure fair allocation of time and memory among all processes while maintaining optimal system performance and enabling processes to complete their tasks as quickly as possible. Many researchers have proposed innovative work to optimise existing process scheduling algorithms, such as Completely Fair Scheduler (CFS). CFS ensures the fair distribution of CPU time among processes in the Linux kernel. However, vanilla CFS faces limitations in multicore environments, where its complexity increases significantly (?). Chen et al. (?) find that vanilla CFS approach maximizes the utilization of processing time but overlooks the contention for lower-level hardware resources. They present a ML-based resource-aware load balancer for the Linux kernel with a low-overhead method for collecting training data. Specifically, they present an ML model based on a multi-layer perceptron (MLP) that imitates the Linux’s CFS load balancer based on the collected training data and an in-kernel implementation of inference on the model. Using eBPF for dynamic tracing, an MLP model replicates CFS decisions, with in-kernel inference for real-time balancing, the model achieves high accuracy and small latency increase, demonstrating effective, low-overhead load management. This scheduler can be extended to bahve with higher efficiency. (?) port this algorithm and model to CUDA and place it in a kernel module to reduce its inference cost.
@@
Goodarzy et al. (?) also question vanilla CFS in its ability for proper allocation of CPU, memory, I/O, and network bandwidth. In response to this, the authors propose SmartOS, an operating system that automatically learns what tasks the user deems to be most important at that time. Based on the learned user preferences, SmartOS adjusts the allocation of system resources such as CPU, memory, I/O, and network bandwidth. It prioritizes the resources for the tasks that the user is currently focused on. The authors demonstrate an implementation of such a learning-based OS in Linux and present evaluation results showing that a reinforcement learning-based approach can rapidly learn and adjust system resources to meet user demands. (?) questions vanilla CFS in its struggles to effectively manage the excessive memory traffic generated by requests for DL-based services. The authors propose a memory-aware fair-share scheduling algorithm designed to reduce the susceptibility of QoS applications to memory-related interference.
Besides process scheduling in the operating systems of individual hosts, ML techniques have been utilized to enhance resource allocation inside computer clusters (e.g., supercomputers). Springborg et al. (?) introduce Chronus, a Python application that collaborates with the Simple Linux Utility for Resource Management (SLURM) scheduler, prevalent in global supercomputers. Chronus executes comprehensive benchmark tests on HPC clusters, varying parameters like core numbers, processor speeds, and hyperthreading. Throughout, it logs performance metrics and energy use for each setup. This data trains a ML model to recognize patterns, aiming to forecast the most energy-efficient configuration for specific jobs and systems. Proven through notable energy savings in benchmarks, this approach holds potential for broader application in HPC systems, promoting eco-friendliness and financial prudence without sacrificing computational power.
From the aforementionend,Compared with traditional schedulers, new scheduling strategies $S$ are often developed because $J(S)$ evolves rapidly and is subject to numerous diverse constraints arising from the complex runtime context. The application of ML to scheduling represents a paradigm shift, enabling more intelligent and dynamic resource allocation as shown in Figure.2. By addressing limitations in traditional algorithms like CFS, ML-based solutions ensure fair resource distribution while optimizing system performance. These advancements, whether in individual hosts or large-scale HPC clusters, demonstrate the potential of ML to transform OS schedulers, balancing fairness, efficiency, and adaptability.
Figure 2: Transition from simple scheduler to ML enhanced scheduler, due to complex running context
Insights and Challenges:
The integration of ML techniques into OS scheduling is driving a paradigm shift, enabling dynamic and intelligent resource allocation that exceeds the limitations of traditional heuristic-based approaches. ML-based schedulers demonstrate remarkable adaptability, capable of optimizing system performance in complex runtime environments by addressing challenges such as workload variability, resource contention, and energy efficiency. Unlike static schedulers, which struggle in current complex environments, these advanced techniques leverage data-driven models to improve prediction accuracy and decision-making. However, these advancements come with significant challenges that must be addressed for widespread adoption. The computational overhead of integrating ML models into real-time scheduling environments poses a critical limitation, as in-kernel inference must balance accuracy with minimal latency to ensure system performance is not compromised. Furthermore, reliance on extensive and high-quality training data for modeling workload patterns and resource utilization introduces complexity in dynamic and heterogeneous environments. Ensuring scalability across large-scale systems and generalizability to unseen scenarios remains difficult, particularly in multicore systems. Trade-offs between competing goals, such as fairness, throughput, latency, and energy efficiency, further complicate optimization, while embedding ML models into existing OS kernels or cluster management systems increases design and implementation complexity. Despite these challenges, the transformative potential of ML-based scheduling is clear, promising more intelligent, adaptable, and efficient resource management solutions for modern computing systems.
4.2 Storage systems
@@
Storage systems, along with their associated OS components, are designed to handle a broad spectrum of applications and fluctuating workloads (?). The storage elements within the OS incorporate a range of heuristic algorithms, ensuring optimal performance and adaptability across diverse workloads. However, these approaches are often limited in their ability to capture the complex, nonlinear relationships between input parameters and performance outcomes.
Predictability is a critical attribute of storage systems, as it ensures consistent and reliable performance across varying workloads and operational conditions. In modern computing environments, where storage systems must support a wide range of applications, from latency-sensitive tasks to high-throughput workloads, predictable behavior is essential for maintaining Quality of Service (QoS) guarantees and meeting Service Level Agreements (SLAs) (?, ?).
@@
With the increasing adoption of Solid-State Drives (SSDs) due to their low cost, high performance, and superior efficiency compared to traditional storage devices, SSDs have become a widely popular choice in storage systems (?). As a result, significant research efforts have been directed toward improving the predictability of SSD performance, aiming to address challenges and ensure reliable operation in diverse application scenarios. Cortez et al. (?) introduce LinnOS, a novel approach to managing SSD performance by incorporating a lightweight neural network within the operating system’s core. This neural network enables real-time inference of SSD performance at an extremely fine-grained level, specifically per-I/O operation, without requiring modifications to the underlying hardware or changes to file systems and applications. By adopting a ’black-box’ perspective towards the device, LinnOS learns and infers the SSD’s behavior to increase predictability, empowering applications to anticipate whether their performance expectations will be met. Compared to traditional methods, LinnOS reduces I/O latency by up to 40% and improves throughput by up to 3× under high-concurrency workloads.